# Generating crossmodal gene expression from cancer histopathology improves multimodal AI predictions

### v3

@@ Acknowledgment
Category AUC ↑ uncertainty ↓ coverage ↑ AUC ↑ uncertainty ↓ coverage ↑ grade uncertainty
Dataset: TCGA - GBMLGG
overall 0.907 0.407 0.981 0.890 0.413 0.981 0.995 0.935
male 0.915 0.439 0.985 0.896 0.444 0.985 0.975 0.920
@@
age $<$ 40 years 0.883 1.000 1.000 0.849 1.000 1.000 0.736 1.000
age from 40 to 60 years 0.928 0.459 1.000 0.921 0.450 1.000 0.804 0.996
age $>$ 60 years 0.909 0.722 0.958 0.906 0.708 0.958 1.000 0.934
magnification 1 0.902 0.195 0.966 0.885 0.172 0.931 1.000 0.652
magnification 2 0.952 0.754 1.000 0.936 0.754 1.000 0.811 1.000
magnification 3 0.852 0.901 1.000 0.836 0.901 1.000 0.795 1.000
alive 0.865 0.990 1.000 0.832 1.000 1.000 0.894 0.881
dead 0.951 0.234 0.973 0.951 0.234 0.973 0.867 1.000
grade II 0.914 0.547 0.974 0.891 0.598 0.974 0.704 0.569
@@
time bin 1 0.898 0.431 0.976 0.882 0.407 0.963 0.984 0.790
time bin 2 0.786 0.538 1.000 0.784 0.513 1.000 1.000 0.817
time bin 3 1.000 0.037 1.000 1.000 0.074 1.000 1.000 0.691
Dataset: TCGA - KIRC
overall 0.778 0.292 0.922 0.773 0.334 0.948 0.850 0.068
male 0.726 0.303 0.909 0.721 0.337 0.929 0.905 0.246
@@
age $<$ 40 years 0.963 0.056 1.000 0.852 0.111 1.000 0.513 0.513
age from 40 to 60 years 0.780 0.393 0.976 0.778 0.389 0.964 1.000 0.894
age $>$ 60 years 0.770 0.295 0.879 0.761 0.318 0.894 0.886 0.521
magnification 1 0.784 0.369 0.980 0.780 0.337 0.961 1.000 0.393
magnification 2 0.780 0.317 0.961 0.774 0.376 0.980 0.867 0.125
magnification 3 0.771 0.337 0.941 0.769 0.363 0.941 0.867 0.495
alive 0.826 0.239 0.931 0.822 0.288 0.951 0.905 0.076
dead 0.661 0.402 0.922 0.656 0.428 0.941 0.880 0.495
grade I 0.222 0.500 0.667 0.111 0.500 0.333 1.000 1.000
@@
time bin 1 0.718 0.366 0.944 0.713 0.384 0.944 0.889 0.565
time bin 2 0.837 0.214 0.883 0.828 0.233 0.900 0.885 0.560
time bin 3 0.848 0.296 1.000 0.835 0.333 1.000 1.000 0.569
time bin 4 0.667 0.111 0.667 0.667 0.167 1.000 1.000 0.513
Table 5: Table of results across various groups for risk estimation. p-values are computed using the Wilcoxon Ranksum Test to compare the scores for grades predicted using real and synthesized transcriptomic data. ↑ indicates the larger the better and ↓ indicates the smaller the better. Equivalent, better or close ($<0.01$) performance with synthesized data is marked in bold.
@@
Category C Index ↑ uncertainty ↓ coverage ↑ C Index ↑ uncertainty ↓ coverage ↑ risk uncertainty
Dataset: TCGA - GBMLGG
overall 0.866 0.446 0.933 0.861 0.438 0.923 0.604 0.652
male 0.808 0.375 0.985 0.813 0.376 0.985 0.956 0.956
female 0.795 0.653 0.921 0.795 0.639 0.921 0.350 0.400
age $<$ 40 years 0.524 0.767 0.977 0.540 0.727 1.000 0.119 0.135
age 40 to 60 years 0.848 0.480 0.973 0.845 0.492 0.973 0.775 0.770
age $>$ 60 years 0.738 0.377 1.000 0.742 0.387 1.000 0.741 0.734
magnification 1 0.855 0.735 0.966 0.840 0.725 1.000 0.646 0.810
magnification 2 0.862 0.484 0.974 0.856 0.486 0.974 0.950 0.975
magnification 3 0.873 0.540 0.973 0.858 0.517 0.973 0.578 0.578
alive - 0.368 0.851 - 0.359 0.851 0.237 0.441
dead 0.614 0.304 0.973 0.633 0.314 1.000 0.563 0.563
grade II 0.823 0.582 0.872 0.823 0.544 0.821 0.058 0.054
grade III 0.590 0.655 0.972 0.526 0.659 0.972 0.901 0.866
grade IV 0.509 0.056 0.966 0.539 0.059 0.966 0.489 0.494
time bin 1 0.827 0.252 0.951 0.822 0.259 0.939 0.942 0.893
time bin 2 - 0.804 0.923 - 0.782 0.923 0.209 0.270
time bin 3 - 0.996 1.000 - 0.984 1.000 0.200 0.233
Dataset: TCGA - KIRC
overall 0.697 0.898 0.882 0.681 0.909 0.928 0.300 0.307
male 0.705 0.909 0.909 0.652 0.921 0.960 0.669 0.654
@@
age $<$ 40 years - 0.955 1.000 - 0.942 1.000 0.275 0.275
age 40 to 60 years 0.612 0.936 0.893 0.591 0.935 0.929 0.266 0.265
age $>$ 60 years 0.672 0.864 0.909 0.681 0.891 0.985 0.781 0.781
magnification 1 0.706 0.932 0.961 0.676 0.947 1.000 0.563 0.588
magnification 2 0.734 0.899 0.902 0.702 0.908 0.961 0.248 0.277
magnification 3 0.694 0.878 0.863 0.676 0.890 0.882 0.408 0.493
alive - 0.934 0.912 - 0.934 0.941 0.083 0.089
dead 0.487 0.845 0.902 0.499 0.874 0.980 0.595 0.597
grade I - 0.976 0.667 - 0.969 1.000 0.275 0.513
@@
time bin 1 0.690 0.922 0.861 0.714 0.938 0.944 0.441 0.474
time bin 2 0.842 0.864 0.983 0.836 0.881 0.983 0.557 0.502
time bin 3 - 0.389 1.000 - 0.389 1.000 0.411 1.000
time bin 4 - 0.775 1.000 - 0.775 1.000 0.275 1.000

### v4

@@ C Index ↑ Wilcoxon test p-value
[…]
overall - 0.907 0.407 0.981 - - 0.890 0.413 0.981 - - 0.995 0.935 -
gender male 0.915 0.439 0.985 0.312 0.953 0.896 0.444 0.985 0.323 0.953 0.975 0.920 0.921
@@
age <40 0.883 1.000 1.000 0.727 0.986 0.849 1.000 1.000 0.720 0.986 0.736 1.000 0.977
40 - 60 0.928 0.459 1.000 0.921 0.450 1.000 0.804 0.996
>60 0.909 0.722 0.958 0.906 0.708 0.958 1.000 0.934
magni- level1 0.902 0.195 0.966 0.617 0.989 0.885 0.172 0.931 0.609 0.977 1.000 0.652 0.884
fication level2 0.952 0.754 1.000 0.936 0.754 1.000 0.811 1.000
level3 0.852 0.901 1.000 0.836 0.901 1.000 0.795 1.000
vital alive 0.865 0.990 1.000 0.612 0.986 0.832 1.000 1.000 0.617 0.986 0.894 0.881 0.941
status dead 0.951 0.234 0.973 0.951 0.234 0.973 0.867 1.000
grade grade II 0.914 0.547 0.974 0.527 0.980 0.891 0.598 0.974 0.544 0.980 0.704 0.569 0.856
@@
time bin bin 1 0.898 0.431 0.976 0.335 0.992 0.882 0.407 0.963 0.331 0.988 0.984 0.790 0.766
bin 2 0.786 0.538 1.000 0.784 0.513 1.000 1.000 0.817
bin 3 1.000 0.037 1.000 1.000 0.074 1.000 1.000 0.691
Category Group Risk Estimation
Real Synthesized p-value ↑
[…]
overall - 0.866 0.446 0.933 - - 0.861 0.438 0.923 - - 0.604 0.652 -
gender male 0.808 0.375 0.985 0.514 0.953 0.813 0.376 0.985 0.507 0.953 0.956 0.956 0.678
female 0.795 0.653 0.921 0.795 0.639 0.921 0.350 0.400
age <40 0.524 0.767 0.977 0.541 0.983 0.540 0.727 1.000 0.535 0.991 0.119 0.135 0.546
40 - 60 0.848 0.480 0.973 0.845 0.492 0.973 0.775 0.770
>60 0.738 0.377 1.000 0.742 0.387 1.000 0.741 0.734
magni- level1 0.855 0.735 0.966 0.586 0.971 0.840 0.725 1.000 0.576 0.982 0.646 0.810 0.787
fication level2 0.862 0.484 0.974 0.856 0.486 0.974 0.950 0.975
level3 0.873 0.540 0.973 0.858 0.517 0.973 0.578 0.578
vital alive - 0.368 0.851 0.336 0.912 - 0.359 0.851 0.337 0.925 0.237 0.441 0.502
status dead 0.614 0.304 0.973 0.633 0.314 1.000 0.563 0.563
grade grade II 0.823 0.582 0.872 0.431 0.937 0.823 0.544 0.824 0.420 0.920 0.058 0.054 0.471
grade III 0.590 0.655 0.972 0.526 0.659 0.972 0.901 0.866
grade IV 0.509 0.056 0.966 0.539 0.059 0.966 0.489 0.494
time bin bin 1 0.827 0.252 0.951 0.684 0.958 0.822 0.259 0.939 0.675 0.954 0.942 0.893 0.465
bin 2 - 0.804 0.923 - 0.782 0.923 0.209 0.270
bin 3 - 0.996 1.000 - 0.984 1.000 0.200 0.233
[…]
Category Group Gradation
Real Synthesized p-value ↑
[…]
overall - 0.863 0.465 0.932 - 0.865 0.465 0.936 - - 0.985 0.991 -
gender male 0.847 0.514 0.972 0.495 0.962 0.847 0.515 0.975 0.494 0.964 0.975 0.968 0.961
female 0.894 0.476 0.953 0.898 0.472 0.953 0.935 0.955
magni- level1 0.861 0.430 0.940 0.478 0.949 0.864 0.435 0.947 0.483 0.956 0.918 0.958 0.938
fication level2 0.865 0.506 0.954 0.867 0.510 0.960 0.931 0.949
level3 0.869 0.497 0.954 0.870 0.503 0.960 1.000 0.905
vital alive 0.835 0.463 0.903 0.478 0.932 0.833 0.465 0.903 0.482 0.938 0.970 1.000 0.947
status dead 0.880 0.493 0.962 0.883 0.499 0.972 0.952 0.894
grade grade IV 0.863 0.465 0.932 - - 0.865 0.465 0.936 - - 0.985 0.991 -
time bin bin 1 0.863 0.465 0.932 - - 0.865 0.465 0.936 - - 0.985 0.991 -
Category Group Risk Estimation
Real Synthesized p-value ↑
[…]
overall - 0.564 0.206 0.971 - - 0.565 0.206 0.971 - - 0.815 0.811 -
gender male 0.653 0.223 0.972 0.205 0.971 0.651 0.222 0.975 0.204 0.973 0.809 0.811 0.870
female 0.430 0.186 0.971 0.432 0.185 0.971 0.931 0.929
magni- level1 0.544 0.208 0.980 0.208 0.971 0.543 0.207 0.987 0.208 0.976 0.718 0.719 0.852
fication level2 0.574 0.209 0.967 0.578 0.209 0.974 0.931 0.937
level3 0.570 0.208 0.967 0.572 0.209 0.967 0.900 0.899
vital alive - 0.183 0.970 0.195 0.964 - 0.184 0.982 0.195 0.974 0.821 0.834 0.853
status dead 0.582 0.207 0.958 0.581 0.206 0.965 0.870 0.873
grade grade IV 0.564 0.206 0.971 - - 0.565 0.206 0.971 - - 0.815 0.811 -
time bin bin 1 0.564 0.206 0.971 - - 0.565 0.206 0.971 - - 0.815 0.811 -
[…]
Category Group Gradation
Real Synthesized p-value ↑
[…]
overall - 0.778 0.292 0.922 - - 0.773 0.334 0.948 - - 0.850 0.068 -
gender male 0.726 0.303 0.909 0.290 0.936 0.721 0.337 0.929 0.337 0.955 0.905 0.246 0.184
@@
age <40 0.863 0.056 1.000 0.248 0.952 0.852 0.111 1.000 0.273 0.953 0.513 0.513 0.643
40 - 60 0.780 0.393 0.976 0.778 0.389 0.964 1.000 0.894
>60 0.770 0.295 0.879 0.761 0.318 0.894 0.886 0.521
magni- level1 0.784 0.369 0.980 0.341 0.961 0.780 0.337 0.961 0.358 0.961 1.000 0.393 0.338
fication level2 0.780 0.317 0.961 0.774 0.376 0.980 0.867 0.125
level3 0.771 0.337 0.941 0.769 0.363 0.941 0.867 0.495
vital alive 0.826 0.239 0.931 0.320 0.926 0.822 0.288 0.951 0.358 0.946 0.905 0.076 0.286
status dead 0.661 0.402 0.922 0.656 0.428 0.941 0.880 0.495
grade grade I 0.222 0.500 0.667 0.356 0.859 0.111 0.500 0.333 0.349 0.779 1.000 1.000 0.625
@@
time bin bin 1 0.718 0.366 0.944 0.247 0.874 0.713 0.384 0.944 0.279 0.961 0.889 0.565 0.552
bin 2 0.837 0.214 0.883 0.828 0.233 0.900 0.885 0.560
bin 3 0.848 0.296 1.000 0.835 0.333 1.000 1.000 0.569
bin 4 0.667 0.111 0.667 0.667 0.167 1.000 1.000 0.513
Category Group Risk Estimation
@@
[…]
overall - 0.697 0.898 0.882 - - 0.681 0.909 0.928 - - 0.300 0.307 -
gender male 0.705 0.909 0.909 0.882 0.834 0.652 0.921 0.960 0.893 0.878 0.669 0.654 0.483
@@
age <40 - 0.955 1.000 0.919 0.934 - 0.942 1.000 0.923 0.971 0.275 0.275 0.440
40 - 60 0.612 0.936 0.893 0.591 0.935 0.929 0.266 0.265
>60 0.672 0.864 0.909 0.681 0.891 0.985 0.781 0.781
magni- level1 0.706 0.932 0.961 0.903 0.908 0.676 0.947 1.000 0.915 0.948 0.563 0.588 0.452
fication level2 0.734 0.899 0.902 0.702 0.908 0.961 0.248 0.277
level3 0.694 0.878 0.863 0.676 0.890 0.882 0.408 0.493
vital alive - 0.934 0.912 0.889 0.907 - 0.934 0.941 0.904 0.961 0.083 0.089 0.343
status dead 0.487 0.845 0.902 0.499 0.874 0.980 0.595 0.597
grade grade I - 0.976 0.667 0.927 0.847 - 0.969 1.000 0.939 0.948 0.275 0.513 0.358
@@
time bin bin 1 0.690 0.922 0.861 0.738 0.961 0.714 0.938 0.944 0.746 0.982 0.441 0.474 0.744
bin 2 0.842 0.864 0.983 0.836 0.881 0.983 0.557 0.502
bin 3 - 0.389 1.000 - 0.389 1.000 0.411 1.000
bin 4 - 0.775 1.000 - 0.775 1.000 0.275 1.000
[…]
Category Group Gradation
Real Synthesized p-value ↑
[…]
overall - 0.828 0.337 0.854 - - 0.821 0.337 0.854 - - 0.941 1.000 -
age 40 - 60 0.739 0.744 0.949 0.489 0.916 0.739 0.744 0.949 0.489 0.916 1.000 1.000 1.000
>60 0.885 0.235 0.882 0.874 0.235 0.882 0.912 1.000
magni- level1 0.881 0.567 0.933 0.519 0.915 0.870 0.544 0.933 0.511 0.915 0.912 0.877 0.959
fication level2 0.810 0.278 0.812 0.810 0.278 0.812 1.000 1.000
level3 0.775 0.711 1.000 0.775 0.711 1.000 1.000 1.000
vital alive 0.811 0.350 0.866 0.300 0.850 0.803 0.350 0.866 0.300 0.850 0.941 1.000 1.000
status dead 0.917 0.250 0.833 0.917 0.250 0.833 1.000 1.000
grade grade I 0.971 1.000 1.000 0.651 0.984 0.971 1.000 1.000 0.651 0.984 1.000 1.000 1.000
grade II 0.805 0.556 1.000 0.768 0.556 1.000 0.893 1.000
grade III 0.823 0.397 0.952 0.823 0.397 0.952 1.000 1.000
time bin bin1 0.761 0.661 0.968 0.470 0.984 0.752 0.651 0.968 0.464 0.984 0.932 0.901 0.950
bin2 0.955 0.278 1.000 0.955 0.278 1.000 1.000 1.000
Category Group Risk Estimation
Real Synthesized p-value ↑
[…]
overall - 0.680 1.000 0.933 - - 0.673 1.000 0.933 - - 0.887 0.992 -
age 40 - 60 - 1.000 0.897 1.000 0.863 - 1.000 0.897 1.000 0.863 0.912 0.980 0.986
>60 0.606 0.999 0.829 0.592 0.999 0.829 0.823 0.992
magni- level1 0.720 1.000 0.900 1.000 0.922 0.680 1.000 0.900 1.000 0.922 0.767 0.971 0.982
fication level2 0.707 1.000 0.933 0.707 1.000 0.933 0.900 0.982
level3 0.547 1.000 0.933 0.547 1.000 0.933 0.918 0.994
vital alive - 1.000 0.936 0.998 0.968 - 1.000 0.936 0.998 0.968 0.897 0.992 0.961
status dead 0.204 0.996 1.000 0.204 0.996 1.000 0.729 0.931
grade grade I - 1.000 0.958 1.000 0.962 - 1.000 0.958 1.000 0.962 0.926 0.992 0.991
grade II 0.778 1.000 1.000 0.746 1.000 1.000 0.805 1.000
grade III 0.503 1.000 0.929 0.509 1.000 0.929 0.936 0.982
time bin bin1 0.714 1.000 0.921 0.996 0.960 0.700 1.000 0.921 0.996 0.960 0.880 0.990 0.970
bin2 - 0.992 1.000 - 0.992 1.000 0.902 0.951
[…]
Category Group Gradation
Real Synthesized p-value ↑
[…]
overall - 0.593 0.821 0.984 - - 0.593 0.818 0.980 - - 1.000 0.914 -
magni- level1 0.595 0.761 0.988 0.805 0.976 0.595 0.761 0.988 0.805 0.976 1.000 1.000 1.000
fication level2 0.601 0.824 0.971 0.600 0.824 0.971 1.000 1.000
level3 0.593 0.831 0.971 0.592 0.831 0.971 1.000 1.000
vital alive 0.575 0.826 0.982 0.717 0.959 0.575 0.823 0.978 0.715 0.957 1.000 0.908 0.954
status dead 0.728 0.608 0.937 0.728 0.608 0.937 1.000 1.000
grade grade I 0.469 0.973 0.960 0.644 0.928 0.465 0.973 0.960 0.643 0.926 0.939 1.000 0.929
grade II 0.454 0.746 0.956 0.454 0.751 0.956 0.952 0.887
grade III 0.887 0.213 0.870 0.889 0.206 0.863 1.000 0.899
time bin bin1 0.496 0.812 0.983 0.841 0.986 0.503 0.809 0.979 0.840 0.984 1.000 0.905 0.953
bin2 0.608 0.870 0.989 0.608 0.870 0.989 1.000 1.000
Category Group Risk Estimation
Real Synthesized p-value ↑
[…]
overall - 0.533 1.000 0.908 - - 0.530 1.000 0.908 - - 0.932 0.918 -
magni- level1 0.393 0.999 0.912 0.999 0.920 0.387 0.999 0.912 0.999 0.920 0.891 0.997 0.950
fication level2 0.594 1.000 0.941 0.592 1.000 0.941 0.975 0.925
level3 0.586 1.000 0.906 0.583 1.000 0.906 0.980 0.927
vital alive - 1.000 0.899 0.582 0.934 - 1.000 0.895 0.583 0.932 0.926 0.993 0.695
status dead 0.509 0.165 0.968 0.510 0.166 0.968 0.996 0.397
grade grade I 0.471 1.000 0.949 0.999 0.921 0.448 1.000 0.960 0.999 0.924 0.871 0.733 0.909
grade II 0.507 1.000 0.933 0.507 1.000 0.933 0.994 0.996
grade III 0.445 0.996 0.879 0.444 0.996 0.879 0.930 0.999
time bin bin1 0.527 1.000 0.915 0.895 0.957 0.527 1.000 0.915 0.895 0.952 0.988 0.999 0.950
bin2 - 0.791 1.000 - 0.790 0.989 0.901 0.901
Table 14: Table of expected deviation from the desired coverage for a given number of samples in the calibration set and conformal error rate $\alpha=0.1$. ncal denotes the calibration set size, and $\epsilon$ denotes the deviation. The less the calibration set size, the greater the expected deviation.
ncal $\epsilon$ ncal $\epsilon$ ncal $\epsilon$ ncal $\epsilon$ ncal $\epsilon$ ncal $\epsilon$
9 0.183 33 0.079 67 0.060 104 0.047 194 0.035 373 0.025
12 0.121 34 0.076 69 0.064 111 0.048 201 0.036 375 0.026
13 0.106 36 0.077 72 0.058 120 0.047 204 0.034 399 0.026
18 0.097 37 0.078 75 0.055 126 0.044 207 0.035 447 0.024
21 0.107 38 0.078 78 0.057 132 0.043 225 0.033 453 0.023
24 0.085 39 0.087 80 0.058 138 0.043 228 0.034 501 0.022
25 0.086 42 0.074 81 0.056 141 0.043 231 0.033 510 0.022
27 0.087 43 0.071 82 0.054 144 0.040 252 0.031 531 0.022
28 0.087 48 0.071 84 0.052 150 0.042 270 0.031 550 0.022
29 0.102 51 0.071 86 0.053 151 0.041 273 0.030 555 0.021
30 0.095 54 0.063 87 0.053 153 0.039 279 0.031 582 0.021
31 0.089 55 0.063 90 0.055 165 0.038 282 0.030 693 0.019
32 0.084 57 0.065 91 0.053 170 0.039 288 0.030 700 0.019
33 0.079 60 0.067 93 0.050 171 0.039 300 0.029 780 0.018
34 0.076 63 0.060 99 0.053 177 0.038 309 0.029 918 0.017
32 0.084 66 0.060 102 0.049 192 0.036 339 0.028 1119 0.015